OpenStudy (anonymous):

The dimensions of a smaller rectangle are 3 ft by 9 ft. The dimensions of a larger rectangle are 5 ft by 15 ft. Find the ratio of the area of the smaller rectangle to the area of the larger rectangle.

OpenStudy (anonymous):

First of all find the area of the smaller rectangle. Do you know the formula for it?

OpenStudy (ghazi):

lets say area of smaller rectangle is A1 and larger A2, now the ratio of areas is given as follows \[\frac{ A _{1}}{ A _{2} }=\frac{ 3*9 }{ 5*15 }=\frac{ 9 }{ 25 }\]

OpenStudy (anonymous):